### v3

# Simpler Optimal Sorting from a Directed Acyclic Graph

Abstract
Fredman proposed in 1976 the following algorithmic problem: Given are a ground set $X$, some partial order $P$ over $X$, and some comparison oracle $O_{L}$ that specifies a linear order $L$ over $X$ that extends $P$. A query to $O_{L}$ has as input distinct $x,x^{\prime}\in X$ and outputs whether $x<_{L}x^{\prime}$ or vice versa. If we denote by $e(P)$ the number of linear orders that extend $P$, then it follows from basic information theory that $\log e(P)$ is a worst-case lower bound on the number of queries needed to output the sorted order of $X$.
Fredman did not specify in what form the partial order is given. Haeupler, Hladík, Iacono, Rozhon, Tarjan, and Tětek (’24) propose to assume as input a directed acyclic graph, $G$, with $m$ edges and $n=|X|$ vertices. Denote by $P_{G}$ the partial order induced by $G$. Their algorithmic performance is measured in running time and the number of queries used, where they use $\Theta(m+n+\log e(P_{G}))$ time and $\Theta(\log e(P_{G}))$ queries to output $X$ in its sorted order. Their algorithm is worst-case optimal in terms of running time and queries, both. Their algorithm combines topological sorting with heapsort, and uses sophisticated data structures (including a recent type of heap with a working-set bound). Their analysis relies upon sophisticated counting arguments using entropy, recursively defined sets defined over the run of their algorithm, and vertices in the graph that they identify as bottlenecks for sorting.
In this paper, we do away with sophistication. We show that when the input is a directed acyclic graph then the problem admits a simple solution using $\Theta(m+n+\log e(P_{G}))$ time and $\Theta(\log e(P_{G}))$ queries. Especially our proofs are much simpler as we avoid the usage of advanced charging arguments and data structures, and instead rely upon two brief observations.
Funding.
Ivor van der Hoog received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 899987. Daniel Rutschmann and Eva Rotenberg received funding from Carlsberg Foundation Young Researcher Fellowship CF21-0302 “Graph Algorithms with Geometric Applications”. Eva Rotenberg additionally receives funding from DFF Grant 2020-2023 (9131-00044B) “Dynamic Network Analysis”, and VILLUM Foundation grant VIL37507 “Efficient Recomputations for Changeful Problems”.
1 Introduction
Sorting is a fundamental problem in computer science. In 1976, Fredman introduced a natural generalisation of sorting, namely sorting under partial information [4]:
Given a ground set $X$ of size $n$, and some partial order $P$ on $X$, and an oracle $O_{L}$ with access to a ground-truth linear order $L$ of $X$, the task is to minimise the number of oracle queries to recover the linear order of $X$. A query takes two input elements from $X$, and outputs their relation in the linear order. Algorithmic efficiency is measured in time, and the number of queries used. Let $e(P)$ denote the number of linear orders that extend the partial order $P$. Then, for any binary (e.g. yes/no) query, $\lceil\log_{2}(e(P))\rceil$ is a worst-case lower bound for the number of queries needed, simply because the sequence of answers should be able to lead to any possible extension of the order.
Previous work.
Fredman [4, TCS’76] shows an exponential-time algorithm for sorting under partial information that uses $\log e(P)+O(n)$ queries. This fails to match the $\Omega(\log e(P))$ lower bound when $\log e(P)$ is sublinear. Kahn and Saks [10, Order’84] prove that there always exists a query which reduces the number of remaining linear extensions by a constant fraction; showing that an $O(\log e(P))$-query algorithm exists. Kahn and Kim [9, STOC’92] are the first to also consider algorithmic running time. They note that Kahn and Saks [10] can preprocess a partial order using exponential time and space, so that given $O_{L}$ it can output the sorted order of $X$ in linear time plus $O(\log e(P))$ queries. Kahn and Kim [9, STOC’92] propose the first polynomial time algorithm that performs $O(\log e(P))$ queries. They do not separate preprocessing the partial order and oracle queries, and use an unspecified polynomial time using the ellipsoid method. Cardinal, Fiorini, Joret, Jungers and Munro [2, STOC’10] (in their full version) preprocess $P$ in $O(n^{2.5})$ time, to output the sorted $X$ in $O(\log e(P)+n)$ time using $O(\log e(P))$ queries. Their runtime poses the question whether query-optimal (sub)quadratic-time algorithms are possible.
For subquadratic algorithms, it becomes relevant how one obtains the partial order. Van der Hoog, Kostityna, Löffler and Speckmann [12, SOCG’19] study the problem in a restricted setting where $P$ is induced by a set of intervals. Here, distinct $x_{i},x_{j}\in X$ are incomparable whenever their intervals $([a_{i},b_{i}],[a_{j},b_{j}])$ intersect (otherwise, $x_{i}\prec_{P}x_{j}$ whenever $b_{i}<a_{j}$). They preprocess the intervals in $O(n\log n)$ time such that, given $O_{L}$, they can produce a pointer to a linked list storing $X$ in its sorted order using $O(\log e(P))$ queries and time. For their queries, they use finger search [1] (i.e., exponential search from a pointer). Finger search has as input a value $x_{i}$, a sorted list $\pi$, and a finger $p\in\pi$ with $p<x_{i}$. It finds the farthest $q$ along $\pi$ for which $q<x_{i}$ in $O(1+\log d_{i})$ time and comparisons. Here, $d_{i}$ denotes the length of the sublist from $p$ to $q$ [1].
Van der Hoog and Rutschmann [8, FOCS’24] assume that $P$ is given as another oracle $O_{P}$ where queries receive distinct $x,x^{\prime}\in X$ and output whether $x\prec_{P}x^{\prime}$. For fixed $c\geq 1$, they preprocess $O_{P}$ using $O(n^{1+\frac{1}{c}})$ time and queries to $O_{P}$. Given $O_{L}$, they produce a pointer to a linked list storing $X$ in its sorted order using $O(c\cdot\log e(P))$ queries and time. Their query algorithm also makes use of finger search. They show matching lower bounds in their setting.
Haeupler, Hladík, Iacono, Rozhon, Tarjan, and Tětek [5, ’24] assume that the partial order $P_{G}$ is induced by a graph $G$ with $m$ edges. Their algorithm can be seen as a combination of topological sort and heapsort. Their algorithm first isolates a collection $B\subset X$ which they call bottlenecks. They then iteratively consider all sources $S$ in $G-B$, remove the source $s\in S$ that is minimum in the linear order $L$, and append $s$ to the output a directed path $\pi$. Before appending $s$ to $\pi$, they find the maximum prefix $B_{s}=\{b\in B\mid b<_{L}s\}$ using finger search where the finger is the head of $B$. They remove $B_{s}$ from $B$, append $B_{s}$ to $\pi$, and then append $s$. To obtain the minimum $s\in S$, they adapt the recently developed heap by Haeupler, Hladík, and Rozhoň, Tarjan and Tětek [6] which has a special working-set bound. For the formal definition, we refer to [5, 6] but intuitively such heaps retrieve the minimum-value $s$ from the heap in time $\log\omega(s)$ instead of $\log n$ where $\omega(s)$ can be thought of as the number of heap elements inserted after $s$. Their algorithm does not separate preprocessing and oracle queries and uses $O(n+m+\log e(P_{G}))$ overall time and $O(\log e(P))$ linear oracle queries. Since reading the input takes at least $O(n+m)$ time, their overall algorithmic runtime is thereby worst-case optimal. The bulk of their analysis is dedicated to charging the algorithmic runtime and query time to the working-set bound of their heap, and to handling the special bottleneck vertices.
Contribution.
We consider the setting from [5], where the input is a directed acyclic graph $G$ over $X$ with $m$ edges and where, unlike [2, 8, 12], one does not separate the algorithmic performance between a pre-processing phase using $G$ and a phase that uses queries to $O_{L}$.
We show that this problem formulation allows for a surprisingly simple solution: Remove a maximum-length directed path $\pi$ from $G$. Iteratively remove an arbitrary source $x_{i}$ from $G$, inserting $x_{i}$ into $\pi$ using finger search where the finger $p_{i}\in\pi$ is the farthest in-neighbour of $x_{i}$ in $G$ along $\pi$ (which we find by simply iterating over all in-neighbors of $x_{i}$). This algorithm relies upon no involved data structures, and instead only uses a finger search tree and lowest common ancestor queries. We use $O(n+m+\log e(P_{G}))$ time and $O(\log e(P_{G}))$ queries. Our proof of correctness is considerably simpler, as it relies upon only one counting argument and one geometric observation.
2 Algorithm
The input is a directed acyclic graph $G$ with vertex set $X=(x_{1},\ldots,x_{n})$. This graph induces a partial order $P_{G}$ where $x_{i}\prec x_{j}$ if and only if there exists a directed path from $x_{i}$ to $x_{j}$ in $G$. The input also contains an oracle $O_{L}$ that specifies a linear order $L$ over $X$. For any distinct $x_{i},x_{j}$, an (oracle) query answers whether $x_{i}<_{L}x_{j}$. The goal is to output $X$ in its sorted order $L$.
We describe our algorithm. For completeness, we compare our algorithm to the one from [5] in Appendix A. Our key observation is that the problem becomes significantly easier when we first extract a maximum-length directed path $\pi$ from $G$. We iteratively remove an arbitrary source $x_{i}$ from $H=G-\pi$ and insert it into $\pi$. Once $H$ is empty, we return $\pi$.
In our analysis, our logarithms are base $2$ and $d^{*}(x_{i})$ denotes the sum of the out-degree and the in-degree of a vertex $x_{i}$ in $G$.
Data structures.
We maintain a path $\pi$ where for all vertices $p,q\in\pi$: $p<_{L}q$ if and only if $p$ precedes $q$ in $\pi$. We store $\pi$ in two leaf-linked binary trees $T_{\pi}$ and $T_{\pi}^{\prime}$. Both trees are subject to finger insertions. $T_{\pi}$ is used for a (finger) search, and $T_{\pi}^{\prime}$ is used for a comparison operation:
FingerInsert$(q_{i},x_{i},T)$. Given vertices $x_{i}\not\in\pi$ and $q_{i}\in\pi$, and a tree $T$ storing $\pi$, insert $x_{i}$ into $\pi$ succeeding $q_{i}$ by replacing the leaf in $T$ storing $q_{i}$ with a node that has $(q_{i},x_{i})$ as its children.
Search$(x_{i},p_{i},O_{L})$. Given a vertex $x_{i}\not\in\pi$ and a vertex $p_{i}\in\pi$ with $p_{i}<_{L}x_{i}$. Return the farthest vertex $q_{i}$ along $\pi$ where $q_{i}<_{L}x_{i}$. Let $d_{i}$ denote the number of vertices on the subpath from $p_{i}$ to $q_{i}$ along $\pi$. We want to use $O(1+\log d_{i})$ time and queries to $O_{L}$. To achieve this, we store $\pi$ in a balanced finger search tree $T_{\pi}$. Many implementations exist [1]. We choose the level-linked (2-4)-tree by Huddleston and Mehlhorn [7], which supports FingerInsert in amortised $O(1)$-time.
Compare$(p,q)$. Given $p,q\in\pi$, return $q$ whenever $q$ precedes $p$ in $O(1)$ time (and $p$ otherwise). We store $\pi$ in a tree $T^{\prime}_{\pi}$ which itself is stored in the dynamic Lowest Common Ancestor (LCA) data structure by Cole, and Hariharan [3]. They maintain $T^{\prime}_{\pi}$ subject to FingerInsert in $O(1)$ time. LCA queries can for any pair of leaves $(u,v)$ give their lowest common ancestor (and whether $u$ is in the left subtree) in $O(1)$ time. This data structure does not support rotating $T^{\prime}_{\pi}$. However, it also does not require $T^{\prime}_{\pi}$ to be balanced. Observe that Compare does not use the oracle $O_{L}$.
Algorithm 1 Sort(directed acyclic graph $G$ over a ground set $X$, Oracle $O_{L}$) time
1:$\pi\leftarrow$ a longest directed path in $G$triangleright $O(n+m)$
2:$T_{\pi}$leftarrow a level-linked (2-4)-tree over $\pi$ [7] $\triangleright$ $O(n)$
3:$T^{\prime}_{\pi}\leftarrow$ a LCA tree over $\pi$ [3] $\triangleright$ $O(n)$
4:$H\leftarrow G-\pi$triangleright $O(n+m)$
5:Compute for each vertex in $H$ its in-degree in $H$triangleright $O(n+m)$
6:$S\leftarrow$ sources in $H$triangleright $O(n)$
7:while  $S\neq\emptyset$ do
8:     Remove an arbitrary vertex $x_{i}$ from $S$triangleright $O(1)$
9:     $p_{i}\leftarrow$ a dummy vertex, which is prepended before the head of $\pi$triangleright $O(1)$
10:     for all in-neighbors $u$ of $x_{i}$ in $G$ do $\triangleright$ $O(d^{*}(x_{i}))$
11:         $p_{i}\leftarrow$ Compare($p_{i}$, $u$)
12:     Remove $x_{i}$ from $H$ and add any new sources in $H$ to $S$triangleright $O(d^{*}(x_{i}))$
13:     $q_{i}\leftarrow$ Search($x_{i}$, $p_{i}$, $O_{L}$) $\triangleright$ $O(1+\log d_{i})$
14:     FingerInsert($q_{i}$, $x_{i}$, $T_{\pi}$$\triangleright$ $O(1)$ amortised
15:     FingerInsert($q_{i}$, $x_{i}$, $T^{\prime}_{\pi}$) $\triangleright$ $O(1)$
16:return the leaves of $T_{\pi}$ in order $\triangleright$ $O(n)$
Algorithm and runtime analysis.
Our algorithm is straightforward. The pseudo code is given by Algorithm 1 where each algorithmic step also indicates the running time. We require linear space. We first extract a longest directed path $\pi$ from $G$, which we store in our two data structures. This takes $O(n+m)$ total time. If $\pi$ has $n-k$ vertices, this leaves us with a graph $H$ with $k$ vertices.
We then iteratively remove an arbitrary source $x_{i}$ from $H$. Since $x_{i}$ is a source in $H$, all vertices $p\in X$ that have a directed path to $x_{i}$ in $G$ must be present in $\pi$. We iterate over all in-neighbours of $x_{i}$ in $G$, and use our Compare operation to find the in-neighbour $p_{i}$ that is furthest along $\pi$. If $x_{i}$ has no in-neighbors, $p_{i}$ is a dummy vertex that precedes the head of $\pi$ instead. Finally, we proceed in a way that is very similar to the algorithm for topological sorting by Knuth [11]: We remove $x_{i}$ from $H$, iterate over all out-neighbors of $x_{i}$ in $H$ and decrement their in-degree, and update the linked list of sources to include the newly found ones. Since we inspect each in- and out-edge of $x_{i}$ only once, this takes $O(m)$ total time.
We then want to find the farthest $q_{i}$ along $\pi$ that precedes $x_{i}$ in $L$. In the special case where $p_{i}$ is the dummy vertex preceding the head $h$ of $\pi$, we use a query to $O_{L}$ to check whether $x_{i}<_{L}h$. If so, we return $q_{i}=p_{i}$. Otherwise, we set $p_{i}=h$ and invoke Search. This returns $q_{i}$ in $O(1+\log d_{i})$ time and queries, where $d_{i}$ is the length of the subpath from $p_{i}$ to $q_{i}$ in $\pi$ ($d_{i}=1$ if $p_{i}=q_{i}$).
Finally, we insert $x_{i}$ into $T_{\pi}$ and $T^{\prime}_{\pi}$ succeeding $q_{i}$ in constant time. We rebalance the tree $T_{\pi}$, which takes amortised $O(1)$ time. Thus, these operations take $O(n)$ total time over all insertions. Therefore, our algorithm uses $O\Bigl{(}n+m+\sum\limits_{x_{i}\in H}(1+\log d_{i})\Bigr{)}$ time and $O\Bigl{(}\sum\limits_{x_{i}\in H}(1+\log d_{i})\Bigr{)}$ queries.
2.1 Proof of optimality
Recall that $e(P_{G})$ denotes the number of linear extensions of $P_{G}$ and let $H$ start out with $k$ vertices. For ease of analysis, we re-index the vertices $X$ so that the path $\pi$ are vertices $(x_{k+1},\ldots,x_{n})$, in order, and $x_{i}$ for $i\in[k]$ is the $i$’th vertex our algorithm inserts into $T_{\pi}$. We prove that our algorithm is tight by showing that $\sum\limits_{i=1}^{k}(1+\log d_{i})=k+\sum\limits_{i=1}^{k}\log d_{i}\in O(\log e(P_{G}))$.
Lemma 1.
Let $G$ be a directed acyclic graph, $P_{G}$ be its induced partial order and $\pi$ be a longest directed path in $G$. If $\pi$ has $n-k$ vertices then $\log e(P_{G})\geq k$.
Proof.
Denote $\ell(x)$ the length of the longest directed path in $G$ from a vertex $x$. Denote for $i\in[n]$ by $L_{i}:=|\{x\in X\mid\ell(x)=i\}|$. For each $i\in[n-k]$, there is one $u\in\pi$ with $\ell(u)=i$ so:
$k=\sum\limits_{i=1}^{n-k}(L_{i}-1).$ (1)
Consider a linear order $L^{\prime}$ that is obtained by first sorting all $x\in X$ by $\ell(x)$ from high to low, and ordering $(u,v)$ with $\ell(u)=\ell(v)$ arbitrarily. The order $L^{\prime}$ must extend $P_{G}$, since any vertex $w$ that has a directed path towards a vertex $u$ must have that $\ell(w)>\ell(u)$. We count the number of distinct $L^{\prime}$ that we obtain in this way to lower bound $e(P_{G})$:
$e(P_{G})\geq\prod_{i=1}^{n-k}L_{i}!\quad\Rightarrow\quad e(P_{G})\geq\prod_{i=1}^{n-k}2^{(L_{i}-1)}=2^{\left(\sum\limits_{i=1}^{n-k}(L_{i}-1)\right)}\quad\Rightarrow\quad e(P_{G})\geq 2^{k}$
Here, the first implication uses that $x!\geq 2^{x-1}$ and the second implication uses Equation 1. ∎
What remains is to show that $\sum\limits_{i=1}^{k}\log d_{i}\in O(\log e(P_{G}))$. To this end, we create a set of intervals:
Definition 1.
Let $\pi^{*}$ be the directed path that Algorithm 1 outputs. For any $i\in[n]$ denote by $\pi^{*}(x_{i})$ the index of $x_{i}$ in $\pi^{*}$. We create an embedding $E$ of $X$ by placing $x_{i}$ at position $\pi^{*}(x_{i})$.
Recall that for $i\in[k]$, $p_{i}$ is the finger from where Algorithm 1 invokes Search with $x_{i}$ as the argument. We create as set $\mathcal{R}$ of $n$ intervals $R_{i}=[a_{i},b_{i}]\subseteq[1,n]$ as follows:
  • If $i>k$ then $[a_{i},b_{i}]:=[\pi^{*}(x_{i}),\pi^{*}(x_{i})]$.
  • Else, $[a_{i},b_{i}]:=[\pi^{*}(p_{i})+1,\pi^{*}(x_{i})]$.
Lemma 2.
Given distinct $x_{i},x_{j}\in X$. If there exists a directed path from $x_{i}$ to $x_{j}$ in $G$ then the intervals $R_{i}=[a_{i},b_{i}]$ and $R_{j}=[a_{j},b_{j}]$ are disjoint with $b_{i}<a_{j}$.
Proof.
We consider three cases.
If $j>k$ and $i>k$ then $x_{i}$ and $x_{j}$ are part of the original longest directed path in $G$. Let without loss of generality $x_{i}$ precede $x_{j}$ on the original longest directed path. Then $\pi^{*}(x_{i})$ precedes $\pi^{*}(x_{j})$ and $R_{i}$ and $R_{j}$ are distinct singletons where $R_{i}$ precedes $R_{j}$.
If $i\leq k$ and $j>k$ then $x_{i}$ was inserted into the path $\pi$ with $x_{j}$ already in $\pi$. Since $\pi^{*}$ is a linear extension of $P_{G}$, it follows that $x_{i}$ was inserted preceding $x_{j}$ in $\pi$. It follows that $x_{i}$ precedes $x_{j}$ in $\pi^{*}$ and thus $R_{i}=[\pi^{*}(p_{i})+1,\pi^{*}(x_{i})$ lies strictly before $R_{j}=\{\pi^{*}(x_{j})\}$.
If $j\leq k$ then the vertex $p_{j}$ must equal or succeed $x_{i}$ in $\pi^{*}$ and thus $b_{i}=\pi^{*}(x_{i})<a_{j}=\pi^{*}(p_{j})$. ∎
The set $\mathcal{R}$ induces an interval order $P_{\mathcal{R}}$ over $X$, which is the partial order $\prec$ where $x_{i}$ and $x_{j}$ are incomparable whenever $R_{i}$ and $R_{j}$ intersect, and where otherwise $x_{i}\prec x_{j}$ whenever $b_{i}<a_{j}$. By Lemma 2, any linear order $L$ that extends $P_{\mathcal{R}}$ must also extend $P_{G}$ and so $e(P_{\mathcal{R}})\leq e(P_{G})$. The number of linear extensions of an interval order are much easier to count. In fact, Cardinal, Fiorini, Joret, Jungers and Munro [2] and Van der Hoog, Kostityna, Löffler and Speckmann [12] already upper bound the number of linear extensions of an interval order. We paraphrase their upper bound (for completeness, we show in Appendix B that this is correct paraphrasing). Let for an interval $R$ with integer endpoints, #$\#R$ denote the number of integers contained in $R$. Then:
Lemma 3 (Lemma 1 in [12] and Lemma 3.2 in [2]).
Let $\mathcal{R}=(R_{1},\ldots R_{n})$ be a set of $n$ intervals in $[1,n]$ with integer endpoints and $P_{\mathcal{R}}$ be its induced partial order. Then $\sum\limits_{i=1}^{n}\log(\#R_{i})\in O(\log e(P_{\mathcal{R}}))$.
We are now ready to prove our main theorem:
Theorem 1.
Given a directed acyclic graph $G$ over $X$, inducing a partial order $P_{G}$, and an oracle $O_{L}$ whose queries specify a linear order $L$ that extends $P_{G}$, there exists an algorithm that uses linear space, $O(n+m+\log e(P_{G}))$ time and $O(\log e(P_{G}))$ oracle queries to output the sorted order of $X$.
Proof.
Our algorithm runs in $O(n+m+k+\sum\limits_{i=1}^{k}\log d_{i})$ time and uses $O(k+\sum\limits_{i=1}^{k}\log d_{i})$ comparisons. By Lemma 1, $k\in O(\log e(P_{G}))$. The set $\mathcal{R}$ of Definition 1 is a set where each interval $R_{i}$ has integer coordinates. By Lemma 2, $e(P_{\mathcal{R}})\leq e(P_{G})$. For $i\leq k$, the size #$\#R_{i}$ must be at least $d_{i}$ since there are per construction at least $d_{i}$ vertices on the subpath from $\pi^{*}(p_{i})$ to $\pi^{*}(q_{j})$ in the embedding $E$. It follows by Lemma 3 that $\sum\limits_{i=1}^{k}\log d_{i}\in O(\log e(P_{\mathcal{R}}))\subseteq O(\log e(P_{G}))$. ∎
Appendix A Comparing Algorithm 1 to the algorithm in [5]
For completeness, we compare the techniques used by our algorithm for sorting from a directed acyclic graph and the algorithm by Haeupler, Hladík, Iacono, Rozhon, Tarjan, and Tětek [5]. In [5], the algorithm does not first extract a maximum-length directed path from $G$.
Instead, they extract bottleneck vertices:
Definition 2 (Section 5 in [5]).
Denote for a vertex $x$ by $\ell(x)$ the length of the longest directed path in $G$ from a vertex $x$. Denote for $i\in[n]$ by $L_{i}:=|\{x\in X\mid\ell(x)=i\}|$. A vertex $x$ is a bottleneck vertex whenever $\ell(x)=i$ and $L_{i}=1$. The set $B$ denotes all bottleneck vertices, sorted by $\ell(x)$.
They prove that whenever $x_{i},x_{j}\in B$ and $\ell(x_{i})<\ell(x_{j})$ then $x_{i}\prec x_{j}$ in the partial order. We observe for the reader that $B$ is some subsequence of all maximum-length paths.
As a data structure, they use an adaption from the min-heap $H$ with working set bound by [6]. The comparison operator of the min-heap is $O_{L}$. For the formal definition of $H$, we refer to [5, 6] but intuitively such heaps retrieve the minimum-value $s$ from the heap in $\log\omega(s)$ time and queries, instead of $\log n$. Here, $\omega(s)$ can be thought of as the number of heap elements inserted after $s$. Their algorithm (Algorithm 2) then works as follows:
Algorithmic description.
We describe the pseudocode and make some remarks upon its workings. Let $Z$ be a graph that starts out as $G$. $B$ starts out as the set of bottleneck vertices in an array. $H$ is the special heap which stores at all times all sources in $Z$ not in $B$. The output $\pi$ is a doubly linked list that starts out empty. They note that the minimum vertex in $Z$, must be either a source in $Z$ or the first vertex of $B$. They iteratively do the following:
First consider the special case that $H$ is empty but $B$ is not. Then the next vertex in $\pi$ must be the first vertex $b$ of $B$. They append $b$ to $\pi$ and remove the out-edges from $b$ from $Z$ and add any new sources in $Z$ (not in $B$) to $H$.
If $H$ is not empty, they remove the minimum vertex $s$ from $H$ (the $s\in H$ where $\forall h\in H$, $h\not<_{T}s$). To find $s$, their heap uses $O(1+\log\omega(s))$ time and queries to $O_{L}$. Before they append $s$ to $\pi$, they first find the maximum prefix $B_{s}$ of $B$ where $\forall b\in B_{s}$: $b<_{T}s$.
  • They use finger search where the finger is the head of $B$.
  • This finds the $d_{s}$ elements in $B_{s}$ in $O(1+\log d_{s})$ queries.
They then append $B_{s}$ and $s$ to $\pi$. They remove $B_{s}\cup\{s\}$ from $Z$, removing all their out-edges from $Z$. Any new source $v$ (with $v\not\in B$) gets added to $H$ using constant time and queries and they recurse. The graph $Z$ is empty if and only if $H$ and $B$ are, at which point they return $\pi$.
Algorithm 2 Topological Heapsort With Search(DAG $G$ over a ground set $X$, Oracle $O_{L}$) [5]
1:$\pi\leftarrow$ an empty list
2:$B\leftarrow$ a set of bottleneck vertices in an array sorted by $\ell(x)$.
3:$Z\leftarrow G$
4:$H\leftarrow$ an empty double pairing min-heap with the working set bound property [6]
5:for all vertices $v$ in $Z$ do
6:     $d(v)\leftarrow$ the in-degree of $v$ in $Z$
7:$H$ adds all sources $s$ in $Z$ where $s\not\in B$
8:while  $H\neq\emptyset$ AND $B\neq\emptyset$ do
9:     if $H=\emptyset$ then
10:         Add the first element of $B$ to $\pi$, remove its out-edges and add all new sources of $Z$ to $H$
11:     else
12:         $s\leftarrow$ the minimum vertex in $H$triangleright Queries to $O_{L}$ using the working set bound
13:         $B_{s}\leftarrow$ maximum prefix of $B$ preceding $s$ in $L$triangleright Finger search over $B$
14:         $B\leftarrow B-B_{s}$
15:         Append $B_{s}$ to $\pi$
16:         Append $s$ to $\pi$
17:         for all edges $(u,v)$ with $u\in B_{s}\cup\{s\}$ do
18:              Decrease $d(v)$ by $1$
19:              if $d(v)=0$ AND $v\not\in B$ then
20:                  add $v$ to the min-heap $H$triangleright $O(1)$ time and queries
21:return $\pi$
A.1 Proof overview and comparison to our analysis
It is straightforward to show that their algorithm uses $O(n+m)$ running time, plus the time spent per source $s$ they remove from $H$. Each time they remove a vertex from $H$ they do two things: they find the minimum $s$ in $H$ (which takes $O(1+\log\omega(s))$ time and queries), and they use exponential search over $B$ (which takes $O(1+\log d_{s})$ time and queries). To prove optimality, they show that:
s $\sum_{s\textnormal{ removed from }H}1\quad+\sum_{s\textnormal{ removed from }H}\log d_{s}\quad+\sum_{s\textnormal{ removed from }H}\log\omega(s)\in O(\log e(P_{G})).$
The first sum they bound in a similar fashion to Lemma 1. Let $k$ denote the number of vertices in $G-B$. They show that $e(P_{G})\in\Omega(2^{k})$. Their bottleneck set does not have the property that for each $i$, there exists one vertex $u\in B$ with $\ell(u)=i$ (and so $k\neq\sum_{i}|L_{i}|-1$). However, they note that for all $i$ for which there is no $u\in B$ with $\ell(u)=i$, $|L_{i}|$ is at least two. This fact can be used to show in a similar fashion that $e(P_{G})\in\Omega(2^{k})$.
For the second sum, their argument essentially considers for each vertex $s$ where $d_{s}>1$, the interval over $B$ over which it performed the finger search. Per construction, the intervals for distinct $s$ and $s^{\prime}$ are disjoint. It follows that there are $d_{s}$ possible number of locations where $s$ could have inserted itself inside this interval without violating the partial order and thus $\prod_{s}d_{s}!\leq e(P_{G})$. The key difference between our arguments, is that we exclusively rely upon finger search. The result of that is that we obtain a set of intervals (in our case over the final ordering $\pi^{*}$) that are not pairwise disjoint. Sorting from a set of intervals is previously studied, and we simply invoke a preexisting lemma from [2, 12].
Finally, we discuss how they upper bound the third sum. This analysis forms the bulk of their paper. We only give a very brief overview and refer to [5]. To relate the working set bound to $\log e(P_{G})$, they recursively define sets over the run of their algorithm:
Definition 3.
Let $S$ be set of vertices. Denote for each iteration $t$ by $S(t)$ the subset of $S$ in the heap at iteration $t$. The greedy sequence $S_{1},S_{2},\ldots$ of a set $S$ as follows: let $t$ be the minimum time such that $|S(t)|$ is maximum and $S^{\prime}=S-S(T)$. If $S^{\prime}$ is empty then the greedy sequence is $\sigma=S(t)$. Otherwise, let $\sigma$ denote the greedy sequence of $S^{\prime}$ then the greedy sequence of $S$ is $\sigma$ with $S(t)$ appended to it.
They use the sizes of $S_{i}$ in the greedy sequence of $G-B$ as an upper bound on the working set size of vertices $s\in S_{i}$ that get removed from the heap. They then relate the term $\sum_{S_{i}\in\sigma}|S_{i}|\log|S_{i}|$ to $\log e(P)$ using entropy of sorted orders. This analysis is vastly different from this paper. We do not use their heap. We instead show that it suffices to iteratively remove an arbitrary source from the current graph, and we exclusively rely upon finger search.
Appendix B Deriving Lemma 3
Lemma 3 is a weaker statement than Lemma 1 from [12] (which itself is paraphrased Lemma 3.2 from [2]). We use Lemma 1 from [12] since its formulation is artithmatically a bit easier to use.
See 3
Proof.
Let $\mathcal{R}$ be set of intervals with integer endpoints. An (open or closed) interval $R_{1}$ precedes an interval $R_{2}$ whenever they are disjoint and each point in $R_{1}$ is to the left of each point in $R_{2}$.
In [12], they define an embedding $\mathcal{I}$ of $\mathcal{R}$ into $[0,1]$ as any set of open intervals $(I_{1},\ldots,I_{n})$ where $I_{i}$ intersects $I_{j}$ if and only if $R_{i}$ intersects $R_{j}$, and $I_{i}$ precedes $I_{j}$ if and only if $R_{i}$ precedes $R_{j}$. They denote by $E(\mathcal{R})$ the set of all embeddings of $\mathcal{R}$ into $[0,1]$ and by $|I_{i}|$ the size of the open interval.
We note that in [12] $H(\mathcal{R})$ denotes the entropy of $\mathcal{R}$ and that $n\cdot H(\mathcal{R})\leq 2\log e(P_{\mathcal{R}})$. Lemma 1 from [12] then states that: [^1] We note that in [12], Lemma 1 uses a maximum which should be a supremum as $E(\mathcal{R})$ has infinite size.
$n\cdot H(\mathcal{R})=\sup_{\mathcal{I}\in E(\mathcal{R})}\left(\sum_{I_{i}\in\mathcal{I}}\log(n|I_{i}|)\right).$
Effectively, the inside of the log scales each open interval in $[0,1]$ to an interval in $[0,n]$ instead, If we denote by $E^{*}(\mathcal{R})$ the set of all embeddings $\mathcal{I}^{*}$ of $\mathcal{R}$ into $[0,n]$, this implies that:
$n\cdot H(\mathcal{R})=\sup_{\mathcal{I}^{*}\in E^{*}(\mathcal{R})}\left(\sum_{I_{i}\in\mathcal{I}^{*}}\log(|I_{i}|)\right).$
The set $\mathcal{R}$ is a set of intervals with integer endpoints in $[1,n]$. We create a set $\mathcal{R}^{\prime}$ of open intervals by adding for each $[a_{i},b_{i}]\in\mathcal{R}$ the interval $(a_{i},b_{i}+1)$. We observe that $\mathcal{R}^{\prime}$ is an embedding $\mathcal{R}^{\prime}$ of $\mathcal{R}$ in $[0,n]$ where each open interval $(a^{\prime}_{i},b^{\prime}_{i})$ has length #$\#[a_{i},b_{i}]$. Since $\mathcal{R}^{\prime}\in E^{*}(\mathcal{R})$ it follows that:
$\sum_{R_{i}\in\mathcal{R}}\log(\#|R_{i}|)=\sum_{R_{i}\in\mathcal{R}}|(a^{\prime}_{i},b^{\prime}_{i})|\leq n\cdot H(\mathcal{R})\leq 2\log e(P_{R})\qed$